OpenStudy (anonymous):

If f(x)=log2(x+4), what is f^-1(3)?

OpenStudy (anonymous):

$$ y = log_2(x+4) $$ Saying \(f^{-1}(3)\) is basically saying \(y = 3\) $$ 3 = log_2(x+4) \\ 2^3 = 2^{(log_2(x+4))} \\ 2^3 = x+4 \\ x = 2^3 - 4 = 8 - 4 = 4 $$ And if we check in the original function then \(f(4)\) is basically: $$ y = log_2(4+4) = log_2(8) = 3 $$

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Sure np =)